# Neutrino masses from new Weinberg-like operators: Phenomenology of TeV scalar multiplets

### v1

@@ 5 Conclusions
In this work we study the possible Weinberg-like operators with new electroweak scalars. We consider scenarios where neutrino masses induced by the SM Weinberg operator are suppressed. These requirements leave us with $2\,(6)$ possibilities with $1\,(2)$ new scalar multiplets up to quintuplet SU(2) representations. We have found that, in some cases, naturally small induced VEVs are generated. Their UV completions lead to new seesaw scenarios, which are studied in a companion paper us:2023two , where we also study the one-loop contributions to neutrino masses and the phenomenology induced by the heavy fermions.
@@ Appendix B Tensor contractions
$\Delta^{\ast}\Phi^{\ast}H\,\Delta=\Delta^{\ast ijk}\Phi^{\ast pqr}H_{i}\Delta_{kqr}\epsilon_{jp}\,,\quad$ $(\Delta^{\ast}\Phi^{\ast}H\,\Delta)^{\prime}=\Delta^{\ast ijk}\Phi^{\ast pqr}H_{s}\,\Delta_{tkr}\epsilon^{st}\epsilon_{ip}\epsilon_{jq}$ (58)
$\Phi^{\ast}\Phi^{\ast}HH=\Phi^{\ast ijk}\Phi^{\ast pqr}H_{i}H_{p}\epsilon_{jq}\epsilon_{kr}\,,\quad$ $\Phi^{\ast}\Phi^{\ast}H\Phi=\Phi^{\ast ijk}\Phi^{\ast pqr}H_{i}\Phi_{kqr}\epsilon_{jp}$
$HHH\Delta=\epsilon^{ij}\epsilon^{kp}\epsilon^{lm}H_{i}H_{k}H_{l}\Delta_{jpm}\,,\quad$ $HH\Phi\Delta=H_{i}H_{j}\Phi_{lmn}\Delta_{pqr}\epsilon^{il}\epsilon^{jp}\epsilon^{mq}\epsilon^{nr}$
$H\Phi\Phi\Delta=H_{i}\Phi_{jkl}\Phi_{mnp}\Delta_{qrs}\epsilon^{ij}\epsilon^{km}\epsilon^{lq}\epsilon^{nr}\epsilon^{ps}\,,\quad$ $\Phi^{\ast}H^{\ast}H\,H=\Phi^{\ast ijk}H^{p}H_{i}H_{j}\epsilon_{kp}$
$\Phi\Phi\Phi\Delta=\Phi_{ijk}\Phi_{pqr}\Phi_{stu}\Delta_{mno}$ $\epsilon^{ip}\epsilon^{jq}\epsilon^{kr}\epsilon^{sm}\epsilon^{tu}\epsilon^{uo}\,.$

### v2

@@ 5 Conclusions
In this work we study the possible Weinberg-like operators with new electroweak scalars. We consider scenarios where neutrino masses induced by the SM Weinberg operator are suppressed. These requirements leave us with $2\,(6)$ possibilities with $1\,(2)$ new scalar multiplets up to quintuplet $SU(2)$ representations. We have found that, in some cases, naturally small induced VEVs are generated. Their UV completions lead to new seesaw scenarios, which are studied in a companion paper Giarnetti:2023osf , where we also study the one-loop contributions to neutrino masses and the phenomenology induced by the heavy fermions.
@@ Appendix B Tensor contractions
$\Delta^{\ast}\Phi^{\ast}H\,\Delta=\Delta^{\ast ijk}\Phi^{\ast pqr}H_{i}\Delta_{kqr}\epsilon_{jp}\,,$ $(\Delta^{\ast}\Phi^{\ast}H\,\Delta)^{\prime}=\Delta^{\ast ijk}\Phi^{\ast pqr}H_{s}\,\Delta_{tkr}\epsilon^{st}\epsilon_{ip}\epsilon_{jq}$ (56)
$\Phi^{\ast}\Phi^{\ast}HH=\Phi^{\ast ijk}\Phi^{\ast pqr}H_{i}H_{p}\epsilon_{jq}\epsilon_{kr}\,,$ $\Phi^{\ast}\Phi^{\ast}H\Phi=\Phi^{\ast ijk}\Phi^{\ast pqr}H_{i}\Phi_{kqr}\epsilon_{jp}$
$HHH\Delta=\epsilon^{ij}\epsilon^{kp}\epsilon^{lm}H_{i}H_{k}H_{l}\Delta_{jpm}\,,$ $HH\Phi\Delta=H_{i}H_{j}\Phi_{lmn}\Delta_{pqr}\epsilon^{il}\epsilon^{jp}\epsilon^{mq}\epsilon^{nr}$
$H\Phi\Phi\Delta=H_{i}\Phi_{jkl}\Phi_{mnp}\Delta_{qrs}\epsilon^{ij}\epsilon^{km}\epsilon^{lq}\epsilon^{nr}\epsilon^{ps}\,,$ $\Phi^{\ast}H^{\ast}H\,H=\Phi^{\ast ijk}H^{p}H_{i}H_{j}\epsilon_{kp}$
$\Phi\Phi\Phi\Delta=\Phi_{ijk}\Phi_{pqr}\Phi_{stu}\Delta_{mno}$ $\epsilon^{ip}\epsilon^{jq}\epsilon^{kr}\epsilon^{sm}\epsilon^{tn}\epsilon^{uo}\,.$